# MACI: Multi-Agent Collaborative Intelligence for Robust Reasoning and Temporal Planning

### v1

@@ Appendix A Experiment Details: Meta-Planning for the Thanksgiving Dinner Task
A.1 Phase 1: Network Construction
A.1.1 Node (Role) Specifications
@@
A.1.2 Edge (Dependency) Specifications
@@
A.2 Phase 2: Agent Assignments
@@
A.2.1 Node (Role) Agent Assignment
@@
A.2.2 Edge (Dependency) Agent Assignment
@@
A.2.3 Common Sense Constraint Analysis (Performed by an LLM)
@@
A.2.4 Common Sense Constraint Analysis and Verification (Human in the Loop)
@@
A.3 Agent Requirements and Assignments
@@
A.4 Monitoring Protocols and Dynamic Adjustments
@@ 3. Coordination Protocol:
A.5 Integrated Workflow Network
@@
A.6 Agent Interaction Specifications
@@
A.7 New Problem Statement Revised with $\mathbf{W^{*}}$
@@
A.8 Experiment #1: Sequential Planner
@@
A.8.1 Results: DeepSeek Wins
@@ A.8.2 Observations on Errors
A.9 Experiment #2: Reactive Planner for Flight Delay
@@
A.9.1 Results: DeepSeek Wins
None of the LLMs cannot react appropriately to this new event without clearing their context buffers. As explained in Appendix A.8.2, this limitation is evident. The key takeaway is that for future runtime frameworks, we must ensure infrastructure support for selectively invalidating stale constraints. If a workflow is already in execution, completed steps and assignments cannot be erased or altered. For example, in a stock-market investment plan, when pertinent news arrives, $\mathcal{MP}$ cannot revert completed nodes or resolved dependencies in $\mathbf{W^{R}}$. For now, we treat the reactive plan as a new plan, given that no steps have been realized in the real world by 10:00 AM.
@@ A.10 Conclusion
Appendix B Validation and Recovery Protocols
$\text{recover}(s_{t},s_{\text{failed}})\rightarrow s_{\text{valid}}$ (16)
B.1 Operations Research Techniques in Validation Protocols
Integration in Multi-Agent Systems
Appendix C $\mathsf{MACI}$ Additional Design Considerations
C.1 Cross-Domain Generalization
C.2 Dynamic Agent Registration and Evolution
C.3 Scalability and Resource Efficiency
C.4 Empirical Evaluation Across Domains
C.5 Challenges and Future Directions
Appendix D Additional Tables and Figures
D.1 Notation
Table 20 presents all symbols used throughout this paper.
Table 20: Symbol Definitions
Symbol Definition Symbol Definition
Basic Sets
$\mathcal{O}$ Planning objectives $\mathcal{P}$ Available people
$\mathcal{M}$ Performance metrics $\mathcal{Q}$ Role qualifications
Workflow Components
$\mathbf{W}$ Workflow network $\mathcal{N}$ Roles (nodes)
$\mathcal{E}$ Dependencies (edges) $\mathbf{A}$ Agent repository
$n_{i}$ Individual role $e_{ij}$ Role dependency
Agent Functions
$f_{\text{role}}$ Role-agent mapping $f_{\text{edge}}$ Edge-agent mapping
Optimization
$A^{*}_{n}$ Selected node agents $A^{*}_{e}$ Selected edge agents
$\mathbf{W^{*}}$ Optimal workflow $V^{*}$ Best validation score
D.2 Example Common Agents
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
Appendix E Traveling Salesman Problem Experiment
E.1 General Problem Specification
  • Inputs: N locations, distance matrix D[N][N]
  • Output: Optimal tour T minimizing total distance
  • Constraints: Each location visited once, return to start
Computational Complexity - Brute Force
For N locations:
  • Number of possible tours = (N-1)!/2
  • Time complexity = O(N!)
  • Space complexity = O(N2)
Solution Methods
  1. 1.
  2. 2.
  3. 3.
E.2 $\mathbf{W^{*}}$: MACI Generated Planner for TSP
Node Components (N)
For TSP with n locations:
$N=\{n_{\text{route}},n_{\text{dist}},n_{\text{valid}}\},\text{ where}$ (17)
    $n_{\text{route}}$: Route generation role
    $n_{\text{dist}}$: Distance calculation role
    $n_{\text{valid}}$: Solution validation role
Edge Dependencies (E)
    $e_{\text{spatial}}$: Distance constraints between locations
    $e_{\text{sequence}}$: Visit order constraints
    $e_{\text{complete}}$: Tour completion requirements
Agent Assignments
Node Agents ($A_{n}$):
  • Route Generation Agent: Generates candidate tours
  • Distance Calculator Agent: Computes tour lengths
  • Solution Validator Agent: Verifies tour validity
Edge Agents ($A_{e}$):
  • Spatial Constraint Agent: Monitors distance feasibility
  • Sequence Monitor Agent: Ensures valid visit order
  • Completion Checker Agent: Verifies tour completeness
Algorithm Selection
Validation Function
E.3 Experiments, From Small to Large N
  1. 1.
    N=5: Establish ground truth via brute force
  2. 2.
    N=10,20,100: Test LLM performance degradation
  3. 3.
    Metrics:
    • Solution quality vs optimal
    • Computation attempts before giving up
    • Error recognition capability
E.3.1 Small Campus Tour (N=5)
Plan an optimal route for campus tour guide visiting 5 key locations:
  • A: Admissions Office (start/end)
  • B: Library
  • C: Student Center
  • D: Science Building
  • E: Sports Complex
Distance Matrix (minutes)
Constraints
  • Tour starts/ends at Admissions (A)
  • Each location visited exactly once
  • Total possible routes: (5-1)!/2 = 12
  • Optimal solution can be verified by hand
3.1.1. $\mathbf{W^{*}}$ Workflow Components
Node Components (N)
  • $n_{\text{route}}$: Generates permutations starting/ending at A.
  • $n_{\text{dist}}$: Computes tour length using distance matrix $D$.
  • $n_{\text{valid}}$: Checks tour validity (start/end at A, no repeats).
Edge Dependencies (E)
  • $e_{\text{spatial}}$: Enforces distance constraints from $D$.
  • $e_{\text{sequence}}$: Ensures visit order consistency.
  • $e_{\text{complete}}$: Validates all 5 locations are visited.
Agent Assignments
  • Node Agents:
    • Route Generation Agent (handles $n_{\text{route}}$)
    • Distance Calculator Agent (handles $n_{\text{dist}}$)
    • Solution Validator Agent (handles $n_{\text{valid}}$)
  • Edge Agents:
    • Spatial Constraint Agent (enforces $e_{\text{spatial}}$)
    • Sequence Monitor Agent (enforces $e_{\text{sequence}}$)
    • Completion Checker Agent (enforces $e_{\text{complete}}$)
Selected Algorithm
Brute-force.
Validation Function
3.1.2. Solution Steps
Step 1: Problem Parsing
  • Input: 5 locations with distance matrix $D$.
  • Initialize node/edge agents and constraints.
Step 2: Route Generation ($n_{\text{route}}$ Agent)
  • Generate all valid permutations: $\frac{(5-1)!}{2}=12$ routes.
  • Example permutations:
Step 3: Distance Calculation ($n_{\text{dist}}$ Agent)
  • Compute total time for each route using $D$.
Step 4: Solution Validation ($n_{\text{valid}}$ Agent)
  • Check all routes for:
    • No duplicate visits.
Step 5: Edge Agent Validation
  • Spatial Constraint Agent: Verify $D_{i,j}$ matches edge weights.
Step 6: Apply Validation Function
  • Assign $V=-\infty$ to invalid routes.
[…]
  • Identify minimal $V=-24$ (i.e., maximal tour length 24 mins).
3.1.3 Solution
Optimal tour time: 242424 minutes, achieved by three routes:
E.3.2 Large Campus Tour (N=10)
Plan an optimal route for a guided tour through 10 locations:
  • Distance Matrix: Asymmetric travel times (minutes)
Planner Best Results Algorithm Iters. Advantages Limitations
Algorithm Selection
3.2.1. ACO Method
Parameters
  • 100 ants, 50 iterations, and $\rho=0.1$ evaporation
  • $\alpha=1$ (pheromone weight), and $\beta=2$ (heuristic weight)
Algorithm
1:  Initialize $\tau_{ij}\leftarrow 1.0$, $\eta_{ij}\leftarrow 1/D_{ij}$
2:  for 50 iterations do
3:     for all 100 ants do
5:        Record tour length $L_{k}$
6:     end for
7:     Evaporate pheromones: $\tau_{ij}\leftarrow 0.9\tau_{ij}$
8:     Deposit pheromones: $\tau_{ij}\leftarrow\tau_{ij}+\sum\frac{10}{L_{k}}$
9:     Track best tour
10:  end for
3.2.2. Performance Comparison
E.4 TSP Experiment Conclusion

### v2

@@ Appendix E Experiment Details: Meta-Planning for the Thanksgiving Dinner Task
E.1 Phase 1: Network Construction
E.1.1 Node (Role) Specifications
@@
E.1.2 Edge (Dependency) Specifications
@@
E.2 Phase 2: Agent Assignments
@@
E.2.1 Node (Role) Agent Assignment
@@
E.2.2 Edge (Dependency) Agent Assignment
@@
E.2.3 Common Sense Constraint Analysis (Performed by an LLM)
@@
E.2.4 Common Sense Constraint Analysis and Verification (Human in the Loop)
@@
E.3 Agent Requirements and Assignments
@@
E.4 Monitoring Protocols and Dynamic Adjustments
@@ 3. Coordination Protocol:
E.5 Integrated Workflow Network
@@
E.6 Agent Interaction Specifications
@@
E.7 New Problem Statement Revised with $\mathbf{W^{*}}$
@@
E.8 Experiment #1: Sequential Planner
@@
E.8.1 Results: DeepSeek Wins
@@ E.8.2 Observations on Errors
E.9 Experiment #2: Reactive Planner for Flight Delay
@@
E.9.1 Results: DeepSeek Wins
None of the LLMs cannot react appropriately to this new event without clearing their context buffers. As explained in Appendix E.8.2, this limitation is evident. The key takeaway is that for future runtime frameworks, we must ensure infrastructure support for selectively invalidating stale constraints. If a workflow is already in execution, completed steps and assignments cannot be erased or altered. For example, in a stock-market investment plan, when pertinent news arrives, $\mathcal{MP}$ cannot revert completed nodes or resolved dependencies in $\mathbf{W^{R}}$. For now, we treat the reactive plan as a new plan, given that no steps have been realized in the real world by 10:00 AM.